Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
SASE1 Status and SASE3 Readiness Technical integration and coordination
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 653     Entry time: 01 Oct 2025, 09:47
Author: Frederik Wolff-Fabris 
Location: General 
Severity: Information 
Subject: Notes XO Integration Meeting 01.10.2025 
General Announcement:
- M. Doering announces on the LIMP-25 for next week: 
1. PLC updates in SA2 and SA3 tunnels (HED postponed to later week); 
2. Karabo deployment happening (no availability); 
3. Timing system update on Tuesday from 09:00 to 16:00 with interruptions during this time; 
4. MKK maintenance on pressurizer for 8C water (e-mail to be sent - please avoid changes on cooling water on Monday); 
5. Magnet cabling in XTD10 starts on Wednesday 08.10.  
- M. Prollius says hot work permits are allowed again.
- I. de Diego says D3 will be testing AIBS / air boxes systems in SA3 experiments tomorrow (access is required).
- Nerea Jardon says please contact EEE if EEE systems need to be operated during the time Karabo is down.

CTRL:
- Preparing for Karabo3 rollout from today.

DA:
- Nothing to report.

EEE:
- Delay on PLC CPU replacement in tunnels: to be done after Karabo is back.
- Timing systems have to be updates - discussed with LAS already and fine with them. 

VAC:
- Everything fine.

XPD:
- Nothing to report.

DET:
- Nothing to report.

ITDM: 
- Long standing issue with SA2 storage system: needs to power cycle system probably tomorrow or Monday, when Karabo is anyway down.
- Update PCs in experiment Hall next week.

XRO:
- Not present.

LAS:
- Nothing to report.

UNSYS:
- Not present.

SPB:
- Nothing to report.

FXE:
- Nothing to report.

MID:
- Nothing to report.

HED:
- Nothing to report.

SCS:
- Nothing to report.

SQS:
- Not present.

SXP:
- Nothing to report.
ELOG V3.1.4-7c3fd00